• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 03:57:38 on August27,2009

    Entry number 286176

    keyword=Left arm cosmics check-out.

    First run after powering back on and redownlaoded the trigger

    S0 looks OK
    VDC is fine besides oddity mentioned in previous entry.

    No signal on HAPPEX detector for now, looks like a pedestal and no TDC. Detector seems to be powered. Maybe someone should check it.

    FIGURE 1